# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

### v2

@@ Abstract
Large language models (LLMs) have facilitated a wide range of applications with distinct service-level objectives (SLOs), from latency-sensitive online tasks like interactive chatbots to throughput-oriented offline workloads like document summarization. The existing deployment model, which dedicates machines to each workload, simplifies SLO management but often leads to poor resource utilization.
This paper introduces HyGen, an interference-aware LLM serving system that enables efficient co-location of online and offline workloads while preserving latency requirements. HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor to estimate batch execution time and an SLO-aware profiler to quantify latency interference, and (2) SLO-aware offline scheduling policies that maximize serving throughput and prevent starvation, without compromising online serving latency. Our evaluation on production workloads shows that HyGen achieves up to 3.87$\times$ overall throughput and 5.84$\times$ offline throughput gains over online and hybrid serving baselines, respectively, while strictly satisfying latency SLOs.
@@ 1 Introduction
Large language models (LLMs) have emerged as transformative tools across diverse domains, handling both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al. (2024a)) and throughput-oriented offline tasks (e.g., document summarization Grand View Research (2023)). Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), while offline tasks prioritize high throughput and resource utilization, often processing large batches with relaxed latency constraints. The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate GPU clusters Sheng et al. (2023); Zhao et al. (2024); Patel et al. (2024); Stojkovic et al. (2024).
However, real-world LLM workloads demonstrate significant temporal variations in request load. Our analysis of production traces shows that in addition to the popular diurnal request arrival patterns, online request rates can vary by up to 3$\times$ within minutes (Section 3). To meet latency requirements under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same GPU cluster. By opportunistically padding online requests with offline requests during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests. However, realizing this opportunity requires addressing several fundamental challenges.
First, LLM services exhibit diverse latency requirements across applications Liu et al. (2024a). Interactive chatbots require consistent response times with both low initial latency (TTFT) and smooth token generation (TBT), while batch processing tasks prioritize throughput over latency. These requirements often manifest in different statistical metrics—from strict P99 latency bounds to mean performance targets—making it difficult to establish unified resource-sharing policies.
@@
Third, co-locating online and offline workloads introduces interference. For instance, offline requests may delay time-sensitive online requests; using large batch sizes to improve offline request throughput can increase the latency of online requests. Effectively managing this interference while ensuring performance guarantees demands meticulous orchestration of resource sharing.
@@
Our evaluation on production workloads shows that HyGen improves serving throughput by 3.87-5.84$\times$ over existing advances Agrawal et al. (2024), while guaranteeing strict SLO compliance. To summarize, this paper makes the following contributions:
  • We investigate the online request patterns in production LLM serving trace, revealing key challenges and opportunities in co-locating online and offline workloads.
  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with performance isolation.
  • We evaluate HyGen using production LLM serving traces to demonstrate HyGen’s effectiveness.
@@ 2 Related Work
LLM Online Serving.
Recent works have improved GPU computation through optimized kernels Dao et al. (2022); Ye et al. (2025), memory efficiency through efficient KV-cache management techniques Lin et al. (2024a); Liu et al. (2024b); Qin et al. (2024), distributed inference through various model parallelism strategies Shoeybi et al. (2019); Pope et al. (2023); Liu et al. (2023), and model compression approaches Xiao et al. (2023); Lin et al. (2024b); Dettmers et al. (2022). Orca Yu et al. (2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching and enables early exits for completed requests. vLLM Kwon et al. (2023) develops PagedAttention for efficient memory management, while Sarathi-Serve Agrawal et al. (2024) proposes chunked prefill to minimize inter-request interference within batches. However, few existing systems address the challenge of co-locating latency-critical online requests with offline workloads.
Throughput-Oriented LLM Inference.
As offline inference gains increasing traction, Huggingface Accelerate Hugging Face (2025), DeepSpeed ZeRO-Inference Aminabadi et al. (2022) and FlexGen Sheng et al. (2023) enable efficient LLM serving on commodity GPUs by offloading strategies. FlexGen aggregates memory from GPU, CPU and disk to maximize serving throughput. S3 Jin et al. (2023) utilizes output sequence length prediction technique, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length.
Co-locating Workloads.
Co-locating latency-sensitive applications with batch applications has been widely adopted to improve resource utilization in datacenters Chen et al. (2019); Zhu & Erez (2016); Zhang et al. (2019); Iorgulescu et al. (2018); Patel & Tiwari (2020); Chen et al. (2023). Prior work has also explored resource multiplexing to serve multiple DNNs Dhakal et al. (2020); Tan et al. (2021); Zhang et al. (2023b), LLMs Duan et al. (2024), or LoRAs Chen et al. (2024); Sheng et al. (2024a); Wu et al. (2024). However, prior efforts have largely overlooked the opportunity to co-locate online and offline inference requests within the same LLM.
@@ 3 Background and Motivation
This section first introduces the core concepts of LLM serving systems (Section 3.1), then illustrates how these characteristics motivate our system designs (Section 3.2).
@@ 3.1 LLM Serving
LLM generation consists of two sequential stages: (1) Prefill Stage: It processes the input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) to avoid regeneration in subsequent tokens, and the (2) Decode Stage: the model generates tokens one at a time in an auto-regressive manner. Each iteration processes only the most recent token while utilizing the KV cache of previous tokens, reducing computational overhead.
Model executions in the prefill stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously. Yet in the decode phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al. (2024), and the total memory access cost grows linearly with the sequence length as each new token’s KV cache is appended.
Serving Workloads.
LLM serving deployment can be categorized into two primary scenarios. The first scenario is online serving, which targets real-time user interactions, such as chatbots, code assistants, and interactive applications OpenAI (2022); Microsoft (2023). This interactive nature often requires short Time to First Token (TTFT)—the latency from receiving a query to generating the first token—as well as the short Time Between Tokens (TBT), i.e., the inter-token generation latency in the decode stage.
The second scenario is offline serving, which prioritizes throughput over latency, measured in queries per second (QPS) or tokens per second (TPS). For example, OpenAI’s Batch API processes requests with relaxed latency requirements (up to 24 hours) at significantly lower costs compared to standard APIs OpenAI (2024). Applications include model benchmarking Hendrycks et al. (2021b); Meta-Team (2024), document processing Chen et al. (2021); Zhang et al. (2023a), data cleaning Narayan et al. (2022); Zheng et al. (2023), and synthetic data generation Allal et al. (2024).
@@ 3.2 Motivation and Challenges
(a) One-hour Request Rates
(b) Two-minute Request Rates
@@
LLM serving systems face a critical resource utilization challenge due to the highly variable nature of their workloads. Figure 1 reports our analysis of Microsoft Azure’s LLM service Patel et al. (2024), showing that request rates can fluctuate dramatically—varying up to 3× within minutes while following broader diurnal patterns. This variability creates an inherent tension in resource provisioning: GPU clusters must be sized to handle peak loads, inevitably leading to resource underutilization during off-peak periods.
This load variability suggests an opportunity to improve GPU utilization by co-locating online requests with offline requests. During periods of low online traffic, the system could opportunistically schedule offline tasks to harvest idle GPU resources. While recent methodologies in simultaneous batching of prefill and decode requests set the premise for dynamic request co-location Agrawal et al. (2024), doing so at scale introduces several fundamental challenges:
@@
    Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements. For example, paid users require strict latency SLOs while free users accept more relaxed guarantees. How to flexiblely specify latency guarantees for online requests when sharing resources?
@@
    Request Interference: Co-locating online and offline workloads introduces performance interference. Large batches of offline request can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests. Worse, batching requests of long inputs with short interactive queries can elongate the latency of all requests in the batch by an order of magnitude Agrawal et al. (2024). How to account for interference in co-locating requests?
@@ 4.1 Overview
Figure 2: HyGen overview. Online requests and offline requests are asynchronously processed. Offline requests are opportunistically scheduled while retaining the batch latency budget.
In this paper, we present HyGen, a novel LLM serving system that enables efficient co-location of online and offline requests while maintaining latency guarantees. Unlike conventional LLM serving systems that segregate online and offline workloads onto separate GPU clusters, HyGen coordinates both types of requests through a two-phase scheduling approach, maximizing resource utilization during off-peak periods while preserving performance isolation.
@@
7:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
8:        if $t_{req}\leq t$ or PHASE ==absent$=$ = ONLINE then
9:           $t\leftarrow t-t_{req}$
10:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
11:        end if
12:     end for
13:     for $r\in R$.prefill $\cup Q$ do
14:        TRY_SCHEDULE:
15:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
@@
17:           $t\leftarrow t-t_{req}$
18:           $c\leftarrow c-l$
19:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
20:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
21:        else
22:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
23:              PERFORM_PREEMPTION($R$, $m$)
24:              goto TRY_SCHEDULE
25:           else
26:              break
27:           end if
28:        end if
29:     end for
30:     return $B$
31:  end function
Figure 2 illustrates the overall design of HyGen. To handle diverse SLOs and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector. This design makes HyGen compatible with various existing scheduling policies for different request types. Following Sarathi-Serve Agrawal et al. (2024), we employ stall-free batching and chunked-prefills, batching prefill and decode requests together for efficient utilization of compute resources and dynamic co-location of requests.
HyGen’s scheduling process occurs in two phases. The first (online) phase handles the online request queue to construct an initial batch, supporting various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al. (2023) and Virtual Token Counter (VTC) Sheng et al. (2024b). If the memory limit is reached and more online requests remain, a priority-based preemption mechanism is employed by prioritizing online requests over offline requests to minimize interference. Preempted offline requests are reclaimed into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
In the second (offline) phase, the remaining capacity is filled with offline requests using a latency predictor to capture the latency impact of request filling. The predictor calculates the latency of the online batch and selects offline requests based on the available batch execution latency budget, i.e., the latency slack without violating the latency requirement of online requests. It predicts the latency cost for decode requests, and calculates a maximum chunked prefill length for prefill requests for the given latency budget. Offline requests are scheduled in order until reaching the resource capacity, based on our SLO-aware scheduling policy. A formalized algorithm of the SLO-aware scheduling process is displayed in Algorithm 1. Appendix A.1 further presents a holistic workflow of two-phase scheduling, including the management of online and offline request queues and implementation of our SLO-aware scheduling algorithm.
Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation. Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining SLOs.
@@ 4.2 Performance Control Mechanisms
Co-locating online and offline requests requires precise control over execution timing and resource allocation. The key challenge lies in accurately estimating the latency impact of scheduling decisions to ensure end-to-end SLO compliance. This section introduces a latency predictor for estimating batch execution latency and an SLO-aware profiler for translating the estimates into scheduling decisions.
@@
We employ linear regression as the prediction model because of its efficiency and effectiveness. Training data for the model is collected by systematically profiling target hardware across diverse batch compositions, varying in the number of requests in different phases, sequence length distributions, and total batch sizes. The linear model enables rapid evaluation of different batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions.
@@
The latency predictor provides accurate latency estimates for filling offline requests. Our SLO-aware profiler leverages a latency budget to ensure SLO compliance in scheduling.
The profiler first analyzes the given combination of workload and SLO to establish viable latency budget ranges. Given that a larger batch size and longer inputs will increase batch latency, the profiler test-runs latency budgets within the range to check their compliance with the given SLO and employs binary search to decide an upper limit that meets the overall SLO for online requests. During deployment, this latency budget is then used as the batch latency limit in the two-phase SLO-aware scheduling process (Section 4.1) to ensure SLO compliance. This profiling enables three key capabilities. First, it determines appropriate latency thresholds that maintain SLO compliance for different workload types. Second, it provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements. Third, it establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.

### v4

@@ Abstract
Large language models (LLMs) have facilitated a wide range of applications with distinct service-level objectives (SLOs), from latency-sensitive online tasks like interactive chatbots to throughput-oriented offline workloads like data synthesis. The existing deployment model, which dedicates machines to each workload, simplifies SLO management but often leads to poor resource utilization.
[…]
@@ 1 Introduction
Large language models (LLMs) have emerged as transformative tools across diverse domains, handling both latency-critical online requests (e.g., chatbot interactions andes ; chatgpt ) and throughput-oriented offline tasks (e.g., document summarization llm-market-report ). Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), while offline tasks prioritize high throughput and resource utilization, often processing large batches with relaxed latency constraints. The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate clusters to avoid interference patel2024splitwise ; sheng2023flexgen ; stojkovic2024dynamollm ; zhao2024blendserve .
[…]
This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same inference engine instance. By opportunistically padding online requests with offline requests during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests. However, realizing this opportunity requires addressing several fundamental challenges.
First, LLM services exhibit diverse latency requirements across applications andes . Interactive chatbots require consistent response times with both low initial latency (TTFT) and smooth token generation (TBT), while batch processing tasks prioritize throughput over latency. These requirements often manifest in different statistical metrics—from strict P99 latency bounds to mean performance targets—making it difficult to establish unified resource-sharing policies.
@@
Third, co-locating online and offline workloads introduces interference. For instance, offline requests may delay time-sensitive online requests; using large batch sizes to improve the throughput of offline requests can increase the latency of online requests. Effectively managing this interference while ensuring latency guarantees demands meticulous orchestration of resource sharing.
@@
[…]
  • Key insights on the feasibility and benefits of co-locating online and offline LLM workloads, derived from systematic characterization of production traces.
  • A statistical latency prediction model that accurately captures the relationship between batch composition and execution latency, accounting for quadratic complexity in prefill and linear scaling in decode phases.
  • A novel scheduling system that dynamically co-locates online and offline workloads, formulated as a constrained optimization problem that maximizes throughput with prefix sharing while preserving strict latency SLOs and providing theoretical guarantees for fairness.
[…]
@@ 2 Related Work
LLM Inference Optimization.
Recent advances in LLM inference through kernel optimization flashattention ; flashinfer , compilation frameworks ma2025intelligen ; wu2025mirage ; zheng2023einnet , and scheduling algorithms lin2024infinite ; lin2024parrot ; sun-etal-2025-disco have substantially improved LLM serving performance. Among these, a particularly important development is predictive scheduling, which aims to estimate request difficulty ding2024hybrid ; ong2024routellm or generation length fu2024efficient ; hu2024inference ; jin2023s ; patke2024queue ; qiu2024power . Contrasting these efforts, HyGen uniquely predicts batch execution time, providing precise control over request interference during co-located workloads. Additionally, the concept of request prefix sharing has been explored by gim2024prompt ; juravsky2024hydragen ; zheng2024sglang , where shared prefixes across requests optimize resource use. HyGen takes this a step further by applying a prefix sharing maximization strategy to opportunistically schedule offline requests, utilizing residual capacity from a primary, SLO-bound online workload. Moreover, HyGen introduces a fairness-aware extension to the standard prefix-sharing maximization (PSM) method, addressing the issue of starvation often encountered in naive designs.
In parallel, the increasing demand for offline LLM inference has driven two key trends. First, platforms like Huggingface Accelerate hugging_face_accelerate , DeepSpeed ZeRO-Inference aminabadi2022deepspeed , and FlexGen sheng2023flexgen enable efficient inference on commodity hardware through memory offloading across GPUs, CPUs, and disk. Second, cloud providers, including OpenAI’s Batch API openai2024batchapi , have launched specialized services optimized for throughput rather than latency, processing millions of requests daily for large-scale data processing tasks.
Workloads Co-location.
In data center environments, co-locating latency-sensitive applications with batch applications has been explored to improve resource utilization chen2023olpart ; chen2019parties ; patel2020clite ; tempo-arixv25 . However, these approaches generally overlook the unique characteristics of LLM inference. In the context of LLM inference, several works have investigated the co-location of various model types and tasks. For instance, Punica chen2024punica , S-LORA sheng2024slora , and dLoRA wu2024dlora batch requests from different LoRAs, while MuxServe duanmuxserve multiplexes resources across multiple LLMs. In contrast, HyGen focuses on batching online and offline requests, addressing distinct optimization problems that are not interchangeable.
@@ 3 Background and Motivation
This section first introduces the background of LLM serving deployment (Section 3.1), then illustrates how these characteristics motivate our system designs (Section 3.2).
@@ 3.1 LLM Serving
Large-scale inference clusters consist of multiple serving instances, with a router intelligently directing incoming requests to the most suitable instances srivatsa2024preble ; iccache-sosp25 . Each instance, which could be based on architectures like vLLM kwon2023efficient or SGLang zheng2024sglang , typically employs iteration-level scheduling yu2022orca and chunked prefill sarathi . This setup enables decode requests to perform an additional decoding step, while prefill requests are limited to a fixed token budget. These two types of requests are processed together within a single iteration, optimizing resource utilization.
LLM serving deployment can be categorized into online serving and offline serving scenarios. Online serving targets real-time user interactions, such as chatbots, code assistants, and interactive applications githubGitHubCopilot ; chatgpt . This interactive nature often requires a short Time to First Token (TTFT) as well as a short Time Between Tokens (TBT). Offline serving prioritizes throughput over latency, measured in queries per second (QPS) or tokens per second (TPS). For example, OpenAI’s Batch API processes requests with relaxed latency requirements (up to 24 hours) at significantly lower costs compared to standard APIs openai2024batchapi . Applications include model benchmarking hendryckstest2021 ; dubey2024llama3herdmodels , document processing chen2021spreadsheetcoder ; zhang-etal-2023-summit , data cleaning narayan2022foundationmodelswrangledata ; zheng2023judgingllmasajudgemtbenchchatbot , and data synthesis huggingfaceCosmopediaCreate .
@@ 3.2 Motivation and Challenges
Figure 2: HyGen Overview. Online and offline requests are processed asynchronously, with offline requests opportunistically scheduled to respect latency budgets.
LLM serving systems face a critical resource utilization challenge due to the highly variable nature of their workloads. Figure 2 reports our analysis of Microsoft Azure’s LLM service patel2024splitwise , showing that request rates can fluctuate dramatically—varying up to 3× within minutes while following broader diurnal patterns. This variability creates an inherent tension in resource provisioning: serving clusters must be sized to handle peak loads, leading to resource underutilization during off-peak periods.
This load variability suggests an opportunity to improve resource utilization by co-locating online requests with offline requests. For each serving instance in the cluster, during periods of low online traffic, it can opportunistically schedule offline tasks to harvest idle resources. While recent methodologies in simultaneous batching of prefill and decode requests set the premise for dynamic request co-location sarathi , doing so at scale introduces several fundamental challenges:
@@
    Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements. For example, paid users require strict latency SLOs while free users accept more relaxed guarantees. How to respect diverse latency requirements in flight?
@@
    Request Interference: Co-locating online and offline workloads introduces performance interference. Large batches of offline requests can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests. Worse, batching requests of long inputs with short interactive queries can elongate the latency of all requests in the batch by an order of magnitude sarathi . How to account for interference in co-locating requests?
@@ 4.1 Overview
HyGen introduces a novel approach to integrate online and offline requests while maintaining strict latency guarantees. As shown in Figure 2, HyGen employs a dual-queue architecture that separates latency-sensitive and throughput-oriented requests. This design accommodates diverse SLOs and variable workloads while remaining compatible with existing scheduling policies within each queue. We note that HyGen functions as an instance-level scheduler, receiving requests from an upstream system-level router (e.g., Preble srivatsa2024preble ). As a result, both the request concurrency and scheduling overhead at each instance are inherently bounded. HyGen’s two-phase scheduling operates as follows (see Appendix A.1 for the asynchronous two-queue workflow and message passing details):
  • The online phase prioritizes latency-sensitive requests, forming an initial batch using established policies such as First-Come-First-Serve (FCFS) kwon2023efficient or fairness request scheduling sheng2024fairness . We introduce a priority-based preemption mechanism that protects online request performance by selectively preempting offline requests. Currently, HyGen preserves execution state for preempted requests, while its architecture supports various preemption mechanisms—including state discarding, preservation, and swapping—as categorized by InferCept abhyankar2024infercept .
  • The offline phase uses our latency predictor to allocate remaining capacity to throughput-oriented requests. This predictor accurately estimates the latency impact of each potential offline request addition, determining either decode request latency costs or maximum chunked prefill lengths that fit within the available latency budget without violating online SLOs.
Algorithm 1 formalizes our scheduling approach, with implementation details and complexity analysis ($O(n)$ where $n$ is the number of requests) provided in Appendices A.1 and A.4. Appendix C presents a novel cluster serving paradigm that addresses the longstanding tradeoff between SLO compliance and resource utilization based on HyGen.
@@
7:   $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)    // predict latency of the decoding request
8:   if $t_{req}\leq t$ or PHASE ==== ONLINE then
9:    // schedule request if it is: 1. online, or 2. offline and enough latency budget left
10:    $t\leftarrow t-t_{req}$
11:    $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
12:  for $r\in R$.prefill $\cup Q$ do
13:   TRY_SCHEDULE:    // try to schedule a prefilling or waiting request
14:   // get the max number of tokens allowed under memory and latency budget
15:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_tokens($t$, $c$, $m$, $r$)
@@
17:    // schedule request
18:    $t\leftarrow t-t_{req}$
19:    $c\leftarrow c-l$
20:    $m\leftarrow m$- GET_NUM_BLOCKS($l$)
21:    $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
22:   else
23:    if PHASE ==== ONLINE and $R\neq B$ then
24:     PERFORM_PREEMPTION($R$, $m$)    // preempt request with lower priority
25:     goto TRY_SCHEDULE    // try to schedule again
26:    else
27:     break
28:  return $B$
@@ 4.2 Performance Control Mechanisms
Co-locating online and offline requests under diverse SLO requirements requires precise control over resource allocation. The key challenge lies in accurately estimating the latency impact of scheduling decisions to ensure SLO compliance. This section introduces a latency predictor for estimating batch execution latency and an SLO-aware profiler for translating the estimates into scheduling decisions.
@@
We employ linear regression as the prediction model because of its efficiency and effectiveness. Training data for the model is collected by systematically profiling target hardware across diverse batch compositions, varying in the number of requests in different phases, sequence length distributions, and total batch sizes. The linear model enables rapid evaluation of varying batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions. Further discussion on the implementation and expandability of the LR predictor can be found in Appendix B.
@@
The latency predictor provides accurate latency estimates for filling offline requests. Our SLO-aware profiler leverages a latency budget to ensure SLO compliance in scheduling. The profiler first analyzes the given combination of workload and SLO to establish viable latency budget ranges. Given that larger batch sizes and longer inputs will increase latency, the profiler test-runs latency budgets within the range to check their compliance with the given SLO and employs binary search to decide an upper limit that meets the overall SLO for online requests. During deployment, this latency budget is used as the batch latency limit in the two-phase scheduling process (Section 4.1) to ensure SLO compliance. This profiling enables three key capabilities: (1) It determines appropriate latency thresholds that maintain SLO compliance for various workloads and limitations (e.g., power constraints patel2024splitwise ). (2) It provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements. (3) It establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.